Molander Corporation is a distributor of a sun umbrella used at resort hotels. Data concerning the next months budget appearEngberg Company installs lawn sod in home yards. The companys most recent monthly contribution format income statement folloRequired: 1. What is the companys degree of operating leverage? 2. Using the degree of operating leverage, estimate the impa
Answer #1

1) Break even unit = Fixed cost/Contribution margin per unit = 9790/26-15 = 890 Unit

Margin of safety ($) = 1040-890*26 = $3900

Margin of safety (%) = (1040-890)/1040 = 14.42%

2a) Degree of operating leverage = Contribution margin/Net income = 89400/70400 = 1.27

2b) Net income increase by = 1.27*28 = 35.56%

2c) Contribution margin income statement

Total Unit
Sales 190720 100%
Variable cost 76288 40%
Contribution margin 114432 60%
Fixed cost 19000
Net income 95432